OpenStudy (anonymous):

1. A music shop has 5 different brands of guitar strings. The 5 different brands sell for $16, $20, $15, $11, and $13. What is the mean absolute deviation of the prices? Round to the nearest cent, if necessary.

OpenStudy (anonymous):

I can help you

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

$12.50 is the mean of the number set

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

you calculate it out on paper and I did no $

OpenStudy (anonymous):

show my plz

OpenStudy (anonymous):

First find the mean number

OpenStudy (anonymous):

okay so you start 16,20,15,11,and 13 then you add them up and 75 about then by 6 and you get $12.50

OpenStudy (anonymous):

divide by 6

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

that is all I did

OpenStudy (anonymous):

can u help with 2 more?

OpenStudy (anonymous):

if you open them in a new quistion

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if so than ya

OpenStudy (anonymous):

\[\frac{ 16+20+15+11+13 }{ 5 }\] sum all and divide them by the quantity of them

OpenStudy (anonymous):

medal please tho first @Alexistheamazing

OpenStudy (anonymous):

You will get 15

OpenStudy (anonymous):

Now you have to graph

OpenStudy (anonymous):

@Alexistheamazing medal first then I will help you

OpenStudy (anonymous):

or mathematically 15 being the center. |16-15|+|20-15|+|15-15|+|11-15|+|13-15| = ?

OpenStudy (anonymous):

thanks let's go now

OpenStudy (anonymous):

The answer is 12

OpenStudy (anonymous):

that divided by 5

OpenStudy (anonymous):

it has to be like this $#.##

OpenStudy (anonymous):

\[\frac{ 12 }{ 5 } = $2.4\]

OpenStudy (anonymous):

This is the final answer

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem.